OpenStudy (aravindg):

prove that IF A+B+C=180, sin^(A/2)+sin^2(B/2)-sin^2(C/2)=1-2cos(A/2)cos(B/2)sin(C/2)

OpenStudy (phi):

This would take me sometime, especially as I am sick today, and a bit fuzzy. I'll give it a shot, but I hope you're not in a hurry.

OpenStudy (aravindg):

:(

OpenStudy (anonymous):

Isn't there square with sin(A/2)?

OpenStudy (aravindg):

YES

OpenStudy (anonymous):

It might be given that A,B,C constitutes a an equilateral triangkle as analysis is saying thjat it will only be true for A=B=C=60.

OpenStudy (phi):

It's definitely true. I started with A+B+C= pi, so A/2 + B/2 + C/2 = 90, and then worked with a=A/2, b=B/2, c= C/2 start with cos(a)cos(b)sin(a) expanded cos(a)cos(b) as 0.5(cos(a-b) + cos(a+b)) It's a bit involved, so I'll type the rest up and post it.

OpenStudy (phi):

Good luck deciphering this